# Cosmological phase-space analysis of $f(G)$-theories of gravity

### v1

@@ I Introduction
Challenging General Relativity has acquired much more attention in view of the most recent cosmological observations that certify the existence of a dark sector 2006IJMPD..15.1753C ; Gruber:2013wua ; delaCruz-Dombriz:2016bqh ; Sahni:2006pa ; Dunsby:2016lkw , fully-unknown, albeit necessary for both clustering and accelerating the universe Bamba:2012cp ; 2001IJMPD..10..213C ; 2003PhRvL..90i1301L . Accordingly, conceptual limitations or experimental drawbacks are becoming gradually more explored Simon2005 , especially in strong gravitational fields 2020A&A…641A.174L ; 2021MNRAS.503.4581L , where gravity breaks down Kiefer:2023bld or a non-trivial mechanism of vacuum energy production is expected Martin:2012bt leading to the cosmological constant problem Luongo:2018lgy ; Belfiglio:2022qai ; Luongo:2023aaq and so on.
@@
To face this problem the renormalization at one loop requires the Einstein-Hilbert action to be supplemented with additional higher-order curvature terms that, however, turn out to be generally not unitary. Analogously, when quantum corrections, for example under the form of string modifications, are included, the effective low-energy gravitational action naturally incorporates higher-order curvature invariants, leading to experimental signature in extremely strong gravity regimes such as cosmological Planck scale, black hole singularities, etc.
@@
However, signatures of this picture may be found even at the level of infrared scales, providing evidences in favor of modifying Einstein’s gravity at all scales.
@@
A natural extension of Einstein’s gravity may lead to $f(R)$ theories of gravity Sotiriou:2008rp ; Nojiri:2008nk ; Nojiri:2008nt , constructed as analytic functions of Ricci scalar $R$, whose prototype is offered by a second order term, $\propto R^{2}$, named scalaron Starobinsky:1980te . Viable $f(R)$ models, fulfilling both cosmological and local gravity constraints, have not really been found so far, and, in general, the matter sector is mainly coupled with such theories, limiting the consistency with observations.
@@
Going through these paradigms, in view of the large number of proposals, one can find theories with a Lagrangian density being a combination of terms Barrow:1988xh ; Barrow:1991hg ; Bueno:2016dol ; Lovelock:1971yv $R_{\mu\nu}R^{\mu\nu}$ and $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$, namely superpositions of the Ricci and Riemann tensors, respectively, that however are jeopardized by unphysical spin-2 ghost instabilities. Analogously, for the sake of completeness, other approaches such as torsion Nesseris:2013jea ; Geng:2011aj ; Paliathanasis:2016vsw ; Duchaniya:2023aeu ; DAgostino:2018ngy or non-metricity Heisenberg:2023lru ; Koussour:2023rly ; Solanki:2022ccf ; Anagnostopoulos:2021ydo can, as well, provide hints on how to heal the aforementioned issue, see e.g. BeltranJimenez:2019esp .
In this respect and quite remarkably, we here single out the study of Gauss-Bonnet (GB) terms Padmanabhan:2013xyr ; Dotti:2007az ; Charmousis:2002rc ; Pozdeeva:2019agu , defined as, $G\equiv R^{2}-4R^{\mu\nu}R_{\mu\nu}+R_{\mu\nu\alpha\beta}R^{\mu\nu\alpha\beta}$, that can significantly improve the above picture.
@@
Over the past decades, various studies have explored this topic, generally revealing that satisfying local gravity constraints remains challenging when the GB term is responsible for dark energy. However, this appears feasible for certain plausible $f(G)$ Lagrangians that yield a cosmological constant Li:2007jm ; Nojiri:2021mxf ; Nojiri:2024nlx ; Myrzakulov:2024sne ; Lohakare:2024ize ; Bamba:2017cjr .
@@
Motivated by the above considerations, we here focus on $f(G)$ theories of gravity and analyze the corresponding stability and phase-space dynamics. To do so, we limit our study to a FLRW background, reconsidering $f(G)$ theories as scalar field-equivalent representation, modifying the corresponding Lagrangian by including a Lagrangian multiplier. After computing the modified Friedmann equations, we formulate the cosmological dynamics by virtue of appropriate dynamical variables, rewritten in order to account for all possible behaviors of the universe, namely enabling the Hubble parameter to change sign, as a consequence of the dynamical prescription. The corresponding autonomous system of dynamical equations has been obtained and the computation of eigenvalues has been found. Analyzing then the phase-space we obtain the existence of attractors and explore the stability of each point under exam. Analogously the stability at asymptotic regime has been found. To do so, we employ a particular version of power-law potential and compare our findings with previous literature. As a matter of fact, we emphasize that the so-constructed model converges to a de Sitter phase
@@ II $f\left(G\right)$-Cosmology
$S[R,G]=\int d^{4}x\sqrt{-g}\left(R+f\left(G\right)\right),$ (2)
@@ II.1 Scalar field equivalent description
Remarkably, introducing a Lagrange multiplier would easily provide an equivalent description in terms of scalar fields. Accordingly, we can show that the geometrodynamical degrees of freedom provided by the nonlinear function $f\left(G\right)$ can be mimicked with a scalar field [^1] A scalar field is commonly associated with interactions and with particles. It is always possible, however, to attribute dynamical properties of a given geometrical model to it, with the great advantage of handling a representation whose stability is easier to compute..
@@
$S_{f\left(G\right)}=\int d^{4}x\sqrt{-g}\left(R+f\left(G\right)-\lambda\left(G-\mathcal{G}\right)\right),$ (3)
@@
that permits to reformulate the Lagrangian by
$\mathcal{L}=R+\left(f\left(G\right)-Gf_{,G}\left(G\right)\right)+f_{,G}\left(G\right)\mathcal{G}.$ (5)
@@
ending up with the equivalent Einstein-Gauss-Bonnet scalar field picture shown as
@@
where we baptize the Hubble function, $H\equiv\frac{1}{N}\frac{\dot{a}}{a}$.
@@ III Cosmological dynamics
$y^{2}=\frac{V\left(\phi\right)}{1+H^{2}},$ (12b)
@@
Further, from Eq. (10) the algebraic constraint follows
@@
By construction, $\eta$ is limited as $\eta^{2}\leq 1$, although the other variables are not compactified. Indeed, for $\eta^{2}=1$, the solution describes a scaling universe, while for $\eta^{2}<1$, the solution describes a de Sitter solution. Nevertheless, when $\eta>0$ the dynamical system describes an expanding universe $\left(H>0\right)$, while for $\eta<0$, the dynamical system describes a collapsing universe $\left(H<0\right)$.
@@
Immediately we remark that,
@@ IV.1 Spatially-flat universe
The equilibrium points with odd numbers describe expanding universes, while the equilibrium points with even numbers describe collapse. Points $A_{5}$, $A_{5}$ describe de Sitter solutions, while the remaining points describe scaling exact solutions.
@@
Equilibrium points $A_{3}$ and $A_{4}$ describe scaling solutions which can describe cosmic acceleration, that is, $q\left(A_{3}\right)=-\frac{4}{3}$ and $q\left(A_{4}\right)=-\frac{4}{3}$. At these points parameter $y$ reaches infinity. The eigenvalues for point $A_{3}$ are $\left\{-4,-\frac{2}{3}\right\}$, while for point $A_{4}$ are derived $\left\{4,\frac{2}{3}\right\}$. Therefore, point $A_{3}$ is always an attractor and point $A_{4}$ is always a source point.
@@ V.1 The role of power-law potential
from where conclude that the point describe an unstable solution. Specifically it is a saddle point for $n>-\frac{1}{\lambda}$, provided that the existence condition requires $\lambda<0$.
@@ VI Final remarks
Accordingly, the scalar field potential is directly related to the function $f(G)$ and its derivatives. Further, within this scalar field formulation, we computed the cosmological field equations and introduced a set of dimensionless variables, quite different from those used in standard approach, i.e., carrying out a global dynamical analysis by enabling the Hubble rate to change its sign.
@@
Indeed, in the power-law $f(G)$-theory, the scaling solution appears absent, and the de Sitter solution is not a future attractor for positive values of the power parameter.
Hence, we considered a generic function $\Gamma(\lambda)$ and, then, we employed the equilibrium points of the dynamical system, described by the exponential potential for each root of the algebraic equation $\Gamma(\lambda_{0})-1=0$. Additionally, they corresponded to the points described by the power-law potential with $x=0$ and an arbitrary $\lambda$. We demonstrated that the form of $\Gamma(\lambda)$ determines the number of distinct roots $\lambda_{0}$ and influences the stability properties of the stationary points.

### v3

@@ I Introduction
Challenging General Relativity has acquired much more attention in view of the most recent cosmological observations that certify the existence of a dark sector 2006IJMPD..15.1753C ; Gruber:2013wua ; delaCruz-Dombriz:2016bqh ; Sahni:2006pa ; Dunsby:2016lkw ; Luongo:2025iqq , fully-unknown, albeit necessary for both clustering and accelerating the universe Bamba:2012cp ; 2001IJMPD..10..213C ; 2003PhRvL..90i1301L . Accordingly, conceptual limitations or experimental drawbacks are becoming gradually more explored Simon2005 , especially in strong gravitational fields 2020A&A…641A.174L ; 2021MNRAS.503.4581L , where gravity breaks down Kiefer:2023bld or a non-trivial mechanism of vacuum energy production is expected Martin:2012bt leading to the cosmological constant problem Luongo:2018lgy ; Belfiglio:2022qai ; Luongo:2023aaq and so on.
@@
To face the problem of renormalization at one loop requires the Einstein-Hilbert action to be supplemented with additional higher-order curvature terms that, mostly turn out to be not unitary. Analogously, when quantum corrections, for example under the form of string modifications, are included, the effective low-energy gravitational action naturally incorporates higher-order curvature invariants, leading to experimental signature in extremely strong gravity regimes such the cosmological Planck scale, black hole singularities, etc.
@@
However, signatures of this picture may be found even at the level of infrared scales, providing evidence in favor of modifying Einstein’s gravity at all scales.
@@
A natural extension of Einstein’s gravity may lead to $f(R)$ theories of gravity Sotiriou:2008rp ; Nojiri:2008nk ; Nojiri:2008nt , constructed as analytic functions of the Ricci scalar $R$, whose prototype is offered by a second order term, $\propto R^{2}$, named scalaron Starobinsky:1980te . Viable $f(R)$ models, fulfilling both cosmological and local gravity constraints, have not really been found so far, and, in general, the matter sector is mainly coupled with such theories, limiting the consistency with observations.
@@
Going through these paradigms, in view of the large number of proposals, one can find theories with a Lagrangian density being a combination of terms Barrow:1988xh ; Barrow:1991hg ; Bueno:2016dol $R_{\mu\nu}R^{\mu\nu}$ and $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$, namely superpositions of the Ricci and Riemann tensors, respectively, that are jeopardized by unphysical spin-2 ghost instabilities. We mention here, for the sake of completeness, other approaches such as torsion Nesseris:2013jea ; Geng:2011aj ; Paliathanasis:2016vsw ; Duchaniya:2023aeu ; DAgostino:2018ngy or non-metricity Heisenberg:2023lru ; Koussour:2023rly ; Solanki:2022ccf ; Anagnostopoulos:2021ydo that can, as well, provide hints on how to facethe aforementioned issue, see e.g. BeltranJimenez:2019esp .
In this respect we single out the study of Gauss-Bonnet (GB) terms Padmanabhan:2013xyr ; Dotti:2007az ; Charmousis:2002rc ; Pozdeeva:2019agu , defined as, $G\equiv R^{2}-4R^{\mu\nu}R_{\mu\nu}+R_{\mu\nu\alpha\beta}R^{\mu\nu\alpha\beta}$, that can significantly improve the above picture. The GB scalar is the first topological scalar provided within the Lovelock theory of gravity Lovelock:1971yv ; Lovelock:1972vz . The latter theory is the second-order alternative to General Relativity gravitational theory in higher-dimensional geometries. However, in the case of a four-dimensional manifold Lovelock’s topological scalars are reduced topopological boundary terms, and general relativity is recovered. Thus, in order to understand the effects of these scalars in low-dimensional geometries, nonlinear terms are introduced in the gravitational action integral.
@@
Over the past decades, various studies have explored this topic, generally revealing that satisfying local gravity constraints DeFelice:2009aj ; Yousaf:2019upm ; Bhatti:2020det remains challenging when the GB term is responsible for dark energy. However, this appears feasible for certain plausible $f(G)$ Lagrangians Nojiri:2005jg that yield a cosmological constant Li:2007jm ; Nojiri:2021mxf ; Nojiri:2024nlx ; Myrzakulov:2024sne ; Lohakare:2024ize ; Bamba:2017cjr ; Nojiri:2010wj ; Nojiri:2017ncd ; Nojiri:2018ouv ; Vernov:2021hxo .
@@
Motivated by the above considerations, we here focus on $f(G)$ theories of gravity and analyze the corresponding stability and phase-space dynamics. To do so, we limit our study to a FLRW background, reconsidering $f(G)$ theories as scalar field-equivalent representation, modifying the corresponding Lagrangian by including a Lagrangian multiplier. After computing the modified Friedmann equations, we formulate the cosmological dynamics by virtue of appropriate dynamical variables, rewritten in order to account for all possible behaviors of the universe, namely enabling the Hubble parameter to change sign, as a consequence of the dynamical behaviour. The corresponding autonomous system of dynamical equations has been obtained and the computation of eigenvalues has been found. Analyzing then the phase-space we obtain the existence of attractors and explore the stability of each point under examination. Following that the stability at asymptotic regime also has been found. To do so, we employ a particular version of power-law potential and compare our findings with previous works. As a matter of fact, we emphasize that the so-constructed model converges to a de Sitter phase
@@ II $f\left(G\right)$-Cosmology
$S[R,G]=\int d^{4}x\sqrt{-g}\left(\frac{R}{2}+f\left(G\right)\right),$ (2)
@@ II.1 Scalar field equivalent description
Remarkably, introducing a Lagrange multiplier would easily provide an equivalent description in terms of scalar fields. Accordingly, we can show that the geometrodynamical degrees of freedom provided by the nonlinear function $f\left(G\right)$ can be mimicked with a scalar field [^1] A scalar field is commonly associated with interactions and particles. It is always possible, however, to attribute dynamical properties of a given geometrical model to it, with the great advantage of handling a representation whose stability is easier to compute..
@@
$S_{f\left(G\right)}=\int d^{4}x\sqrt{-g}\left(\frac{R}{2}+f\left(G\right)-\lambda\left(G-\mathcal{G}\right)\right),$ (3)
@@
that permits to reformulate the Lagrangian as
$\mathcal{L}=\frac{R}{2}+\left(f\left(G\right)-Gf_{,G}\left(G\right)\right)+f_{,G}\left(G\right)\mathcal{G}.$ (5)
@@
ending up with the equivalent Einstein-Gauss-Bonnet scalar field picture
@@
where we consider the Hubble function to be, $H\equiv\frac{1}{N}\frac{\dot{a}}{a}$.
@@ III Cosmological dynamics
$y=\frac{V\left(\phi\right)}{1+H^{2}},$ (12b)
@@
Also, from Eq. (10) the algebraic constraint becomes
@@
By construction, $\eta$ is limited as $\eta^{2}\leq 1$, although the other variables are not compactified. Indeed, for $\eta^{2}=1$, the solution describes a scaling universe, while for $\eta^{2}<1$, the solution describes a de Sitter solution. When $\eta>0$ the dynamical system describes an expanding universe $\left(H>0\right)$, while for $\eta<0$, the dynamical system describes a collapsing universe $\left(H<0\right)$.
@@
Immediately we notice that,
@@ IV.1 Spatially-flat universe
The equilibrium points with odd numbers describe expanding universes, while the equilibrium points with even numbers describe collapse. Points $A_{5}$, $A_{6}$ describe de Sitter solutions, while the remaining points describe scaling exact solutions.
@@
Equilibrium points $A_{3}$ and $A_{4}$ describe scaling solutions which can describe cosmic acceleration, that is, $q\left(A_{3}\right)=-\frac{4}{3}$ and $q\left(A_{4}\right)=-\frac{4}{3}$. At these points parameter $y$ reaches infinity. The eigenvalues for point $A_{3}$ are $\left\{-4,-\frac{2}{3}\right\}$, while for point $A_{4}$ are $\left\{4,\frac{2}{3}\right\}$. Therefore, point $A_{3}$ is always an attractor and point $A_{4}$ is always a source point.
@@ IV.2.1 Analysis at infinity
In term of dynamics the curvature component can be seen equivalently in the same form of a spatially flat geometry induced with a indeal gas with constant equation of state $p_{m}=-\frac{1}{3}\rho_{m.}$. By using this analogue, we can compare our results with the matter dominated investigated in Goheer:2009qh . Nevertheless this is an asymptotic solution, near a specific value of the Gauss-Bonnet scalar because $x=0$, that is, the function (20 ) is described by a polynomial similar to that determined in Goheer:2009qh .
@@ V.1 The role of power-law potential
from where conclude that the point describes an unstable solution. Specifically it is a saddle point for $n>-\frac{1}{\lambda}$, provided that the existence condition requires $\lambda<0$.
@@ V.2 Asymptotic analysis at Infinity
V.3 Discussion for a general potential
Consider now the generic nonlinear potential $V\left(\phi\right)$, that is, a generic nonlinear function $f\left(G\right)$. Then the equilibrium points solve the algebraic equation (17), that is,
$\frac{\lambda}{8}x\left(\Gamma\left(\lambda\right)-1\right)=0$
that is, $x=0$ or $\lambda=\lambda_{0}$ such that $\lambda_{0}\left(\Gamma\left(\lambda_{0}\right)-1\right)=0$, where $\Gamma\left(\lambda\right)$ is constructed by the potential.
When $x=0$, it follows from before that $\lambda$ does not constrained by the rest equations of the dynamical system, that is, the admitted equilibrium points are those described by the power-law potential.
On the other hand, for $\lambda=\lambda_{0}$, then from the rest of the equations variables $x,\eta$ are constrained and provide families of equilibrium points as given by the exponential potential, for each value of $\lambda_{0}$.
Therefore, the definition of an arbitrary potential function will not lead to the derivation of new families of solutions. Hence, the study of the exponential and the power-law potentials is sufficient to construct all the possible asymptotic solutions.
However, what is important to mention is that the stability properties of the equilibrium points will not remain the same. Because the function $\Gamma\left(\lambda\right)$ and, its derivative play an important role in the linearization of the dynamical system around the equilibrium points.
@@ VI Final remarks
The scalar field potential is directly related to the function $f(G)$ and its derivatives. Further, within this scalar field formulation, we computed the cosmological field equations and introduced a set of dimensionless variables, quite different from those used in standard approach, i.e., carrying out a global dynamical analysis by enabling the Hubble rate to change its sign.
@@
Indeed, in the power-law $f(G)$-theory, the scaling solution appears absent, and the de Sitter solution is a future attractor for positive values of the power parameter and for specific set of initial conditions. That means that while the $f(G)$-theory can provide an explanation to the cosmic acceleration, the majority of the trajectories of the solutions lead to attractors which can not describe the present cosmic acceleration. Nevertheless, unstable equlibrium points which describe cosmic acceleration can be related to the early-time inflation.